NAME MATH 164, $6.2 WORKSHEET SPRING 2020 PERRY ALIADO Calculate the mean, standard deviation, and probability for the follow please answer all 3 circled questions
Answer #1

7 Calculate mean standard deviation, and albe probability for the following benomial experiments with the following numericaland the raciance of Binomial distribution is Variance = npq. = 8* 0.836 x 0.164 variance = 1.097 and standard deviation is roAchinity: Picking domi onnity: Picking dominoee luging Double sin domino set Number of times coe pick a domino, then put it= 6! 0.0625 X 10.5)4 (6-2)!2! 2 = 96-45x400625 x 0. OLDAL E4 X2X1 wurde ind = 15 x 0.0625 %0.316de = 0.29625 1 PCx=2] = 0.30and variance = npq 02= 9x0.472 X 0.528 72 - 2.942944 DE 12.242944 = 7:49764615 :18.0.21.50 Probability for the Binomial dig h